OpenStudy (quickshadow):

Two equations are given below: a − 3b = 16 a = b − 2 What is the solution to the set of equations in the form (a, b)?

OpenStudy (lyss):

Is this sort of like x-3y=16, x=y-2 ?

OpenStudy (will.h):

we can solve using substitution method substitute the 2nd equation into the 1st (b - 2) - 3b = 16 solve for b -2b - 2 = 16 -2b = 18 b = -9 now solve for a in the 2nd equation by substituting the value of b a = -9 -2 a = -11 Therefor the answer is (-11,-9)